# Deterministic Dynamic Edge-Colouring

### v1

@@ 5 Maintaining a hierarchy of $t$-splitters
In this section, we will show how to build a hierarchy of dynamic $t$-splitters in order to split the degree all the way down to $O(\operatorname{\text{{\rm poly}}}(\varepsilon^{-1},\log n))$ while retaining an update time of $2^{\tilde{O}_{\log\varepsilon^{-1}}(\sqrt{\log n})}$. Thus we will show Theorem 4, which we restate below:
Theorem 30 (Identical to Theorem 4).
Let $G$ be a dynamic graph, and let $01>\varepsilon>$ 0 and an upper bound $\Delta_{\max}\geq\tfrac{10^{7}\log^{5}n}{\varepsilon^{3}}$ on the maximum degree of $G$ through the entire update sequence be given. Set $t_{1}=\lfloor 2^{10\sqrt{\log n}}\rfloor$ and $t_{2}=\lfloor\tfrac{\log n}{\varepsilon}\rfloor$. Then there exists a dynamic algorithm maintaining parameters $h,i,j$ with $h=i+j\leq 4\sqrt{\log n}$ and a set of graphs $\mathcal{G}_{h}$ satisfying that:
@@
  2. 2.
    $\hat{\Delta}_{i+j}\leq(1+\tfrac{\varepsilon}{16})\tfrac{\Delta_{\max}}{t_{1}^{i}t_{2}^{j}}$
@@ 5.1 The hierarchy
We first describe the hierarchy in words. The hierarchy will consist of two types of $t$-splitters. The first type of $t$-splitter, we will use, has $t_{1}=\lfloor 2^{10\sqrt{\log n}}\rfloor$, and the second type has $t_{2}=\lfloor\tfrac{\log n}{\varepsilon}\rfloor$. Given any graph $G$, any $\Delta_{\max}$, and any $10<\varepsilon<$ 1, we will select some $h$. The hierarchy then consists of a sequence of sets of graphs $\{\mathcal{G}_{i}\}_{i=0}^{h}$ and a sequence of values $\{\hat{\Delta}_{i}\}_{i=0}^{h}$ such that for all $0\leq i\leq h$ and all $H\in\mathcal{G}_{i}$, we have that $\max_{v\in V(H)}d(v)\leq\hat{\Delta}_{i}$ i.e. the sequence of $\hat{\Delta}_{i}$ provides a sequence of upper bounds on the max degrees.
@@
We note already here that since the algorithm from Theorem 3 is not history independent and so ’running’ it on a static graph is ill-defined. We will ignore this in the following informal description, and then provide a proper definition afterwards. We will initialise $\mathcal{G}_{0}=\{G\}$ and $\hat{\Delta}_{0}=\Delta_{\max}$, and proceed as follows: given $\mathcal{G}_{i}$ and $\hat{\Delta}_{i}$, we let $\mathcal{G}_{i+1}$ be the set of graphs obtained in the following manner. If $\Delta_{\max}\geq 2^{10i\sqrt{\log n}}$, we will run the algorithm from Theorem 3 with the parameter $\eta=\tfrac{\varepsilon^{\prime 2}}{\log n}$ and $t=t_{1}$ on all $H\in\mathcal{G}_{i}$. For each $H$, every colour-class $\kappa$ induces a subgraph $H_{\kappa}$ of $H$. For each choice of $H$ and $\kappa$, we include $H_{\kappa}$ in $\mathcal{G}_{i+1}$ after removing all degree 00 vertices. Finally, by Theorem 3, we can choose $\hat{\Delta}_{i+1}=(1+\eta)\tfrac{\hat{\Delta}_{i}}{t}$. Otherwise we set $t=t_{2}$ and repeat the above construction. We will pick $h$ such that the construction terminates with $\hat{\Delta}_{h}=O(\operatorname{\text{{\rm poly}}}(\varepsilon^{-1},\log n))$, but still large enough for any application of Theorem 3 to go through.
@@
The idea is that the edge sets of the graphs in $\mathcal{G}_{h}$ form a partition of the edges of $G$ such that each partition induces a low-degree subgraph of $G$. Since the $t$-splitters used are fully-dynamic, we can maintain the above hierarchy dynamically as well (see Algorithms 7 8 9). To do so, we simply build the $t$-splitters as specified, and whenever an edge is inserted, deleted or recoloured, we remove it from the graphs it previously belonged to (if any) and instead insert it into the graphs specified by its new colour (if any). The order of operations does not matter too much, as long as we guarantee that any $t$-splitter algorithm that is called is allowed to finish and that the $\hat{\Delta}_{i}$’s remain valid upper-bounds through-out the update sequence. In Algorithms 7 8 9 we will explicitly fix the order of operations.
@@
Next, we introduce some basic notation in order to make the above description formal. First of all, we define a $(G,\Delta_{\max},\varepsilon)$-valid-hierarchy. This is any hierarchy obtained by splitting via any of $t$-splitter satisifying Invariant 20, and not necessarily only via the dynamic $t$-splitter from Theorem 3. This deals with the technicality that the data structure from Theorem 3 might not be history-independent, and so we cannot dynamically hope to maintain a very strictly defined hierarchy.
Definition 31.
@@
  • $\mathcal{G}_{i+1}:=\{\hat{H}:\exists H\in\mathcal{G}_{i},\kappa\in[t^{\prime}]\text{ with }H[\kappa]=\hat{H}\text{ under }c_{H}\}$.
  • For all $i$: $\hat{\Delta}_{i+1}\leq(1+\tfrac{\varepsilon}{128\log n})\tfrac{\hat{\Delta}_{i}}{t^{\prime}}$
@@
Remark 32.
We remark that the choice of the value $2^{(10+(i-1))\log n}$ is not extremely precise. One could have equally well chosen $2^{(9+(i-1))\log n}$ or $20\cdot{}2^{(10+(i-1))\log n}$. One then need only change the constant in Algorithms 7 9 8, in Definition 31, and in the subsequent proofs mutatis mutandi. In particular, one can always discretize the values and slightly over- or under-estimate all parameters like $\Delta_{\max},\varepsilon$ or $2^{(10+(i-1))\log n}$ to ensure that one does not need access to real RAM in order to do fast comparisons.
In particular, any hierarchy that uses a valid instantiation of Theorem 3 for every $t$-splitter yields a $(G,\Delta_{\max},\varepsilon)$-valid-hierarchy; regardless of the insertion order. We will also observe that for all $i$, the edge sets of the graphs in $\mathcal{G}_{i}$ form a valid partition of $E(G)$. We will let a vector $\chi(e)$ describe the hierarchy-profile of an edge:
Definition 33.
@@
For any $i\leq h$ and any vector $v\in\mathbb{Z}^{h}$, we let the $i^{\text{th}}$ initial segment of $v$ be the vector $v^{(i)}\in\mathbb{Z}^{i}$ whose first $i$ coordinates agree with the first $i$ coordinates of $v$. For any vector $v\in\mathbb{Z}^{i}$ and constant $k\in\mathbb{Z}$, we also let $[v^{(i)},k]$ be the vector in $\mathbb{Z}^{i+1}$ with $i^{th}$ initial segment equal to $v$ and $[v,k]_{i+1}=k$ We can index the graphs in the hierarchy by providing an index set $J$ containing every initial segment of the hierarchy-profile of every edge $e\in G$. Here, we let $G_{\emptyset}=G$, and we let $G_{j}$ for any $j\in J$ be the graph containing exactly the edges $e$ for which $j$ is an initial segment of $\chi(e)$. We let $|j|:=dim(j)$ denote the dimension of the vector $j$. Also $|j|$ is the level of $G_{j}$.
@@ 5.2 An algorithm for maintaining a hierarchy
The algorithm for dynamically maintaining the hierarchy is quite straight-forward. First of all, one initializes values $t(i)$ and $\hat{\Delta}_{i}$ so that they conform to the constraints of Definition 31 (see Algorithm 7). This only has to be done once.
The hierarchy is then maintained by repeatedly applying Theorem 3 to further split graphs into ones with lower degree. First one performs the update on level 00, which then propagates the update along with any extra recolourings to the next level of the hierarchy. One has to be a bit careful with determining the order of deletions and insertions, in order to ensure that the $\hat{\Delta}_{i}$’s remain valid upper-bounds through-out the update sequence. Therefore, the algorithm first processes an update in the top level, deletes all edges that were recoloured from the old graphs they were put into, and then inserts any new edges into the further recursive structures. This ensures that $\hat{\Delta}_{i}$’s remain valid upper-bounds. Finally, we note that the hierarchy consists of many graphs, which can either be constructed in a pre-processing phase or simply allocated online by only inserting vertices into a graph once the vertex actually has degree at least one.
Next, we give pseudo-code for the above dynamic algorithm. If an edge $e$ has just been recoloured, we let $c(e)_{\text{old}}$ respectively $\chi(xy)_{\text{old}}$ denote the colours of $e$ just before the recolouring. The hierarchy is maintained via the following set of recursive procedures. The first procedure shows how to initialize the parameters needed for the following procedures. We note already now that to actually perform an insertion or a deletion, one must call the respective recursive procedure with the correct parameter setting (see Remark 34):
@@
  $\eta=\lfloor\tfrac{\varepsilon}{128\log n}\rfloor$
@@
        $\hat{\Delta}_{i_{1}+i_{2}}=(1+\eta)^{i_{1}+i_{2}}\tfrac{\Delta_{\max}}{t_{1}^{i_{1}}t_{2}^{i_{2}}}$
@@
        $\hat{\Delta}_{i_{1}+i_{2}}=(1+\eta)^{i_{1}+i_{2}}\tfrac{\Delta_{\max}}{t_{1}^{i_{1}}t_{2}^{i_{2}}}$
@@
Algorithm 7 Initialize-Parameters$(\Delta_{\max},\varepsilon)$
  insert(uv) into $G_{j}$ using Theorem 3 with $(\Delta_{\max},t,\varepsilon^{\prime})=(\hat{\Delta}_{|j|},t(|j|),\eta)$
@@
     for $k=h$ to $h-|j|$ do
        Delete-Hierarchy$(xy,\chi(xy)^{(k)}_{\text{old}},\chi(xy)_{\text{old}})$
     Insert-Hierarchy$(xy,[j,c_{G_{j}}(xy)])$
  Insert-Hierarchy$(uv,[j,c_{G_{j}}(uv)])$
Algorithm 8 Insert-Hierarchy$(uv,j)$
  delete(uv) from $G_{j}$ using Theorem 3 with $(\Delta_{\max},t,\varepsilon^{\prime})=(\hat{\Delta}_{|j|},t(|j|),\eta)$
  if $|v|\geq|j|+1$ then
     Delete-Hierarchy$(uv,v^{(|j|+1)},v)$
  for all recoloured edges $xy\neq uv$ in $G_{j}$ do
     for $k=h$ to $h-|j|$ do
        Delete-Hierarchy$(xy,\chi(xy)^{(k)}_{\text{old}},\chi(xy)_{\text{old}})$
     Insert-Hierarchy$(xy,[j,c_{G_{j}}(xy)])$
Algorithm 9 Delete-Hierarchy$(uv,j,v)$
Remark 34.
[…]
5.3 Correctness
In this section, we show that the algorithms from above actually maintain a $(G,\Delta_{\max},\varepsilon)$-valid-hierarchy with parameters that give the guarantees stated in Theorem 4. We first show that the algorithm maintains a $(G,\Delta_{\max},\varepsilon)$-valid-hierarchy:
Lemma 35.
Algorithms 7 8 9 implement a $(G,\Delta_{\max},\varepsilon)$-valid-hierarchy.
Proof.
@@ 5.3 Correctness
The hierarchy is also in accordance with the third, fifth, and sixth items. Indeed, if at all times the conditions of Theorem 3 is met, then it follows by induction and the theorem itself that the computed colourings are $t(i)$-splitting for all $i$, and furthermore satisfy the fifth and sixth items.
To check that this is the case, we check that the upper bounds on the degrees $\hat{\Delta}_{i}$ remain valid at all times during an update. This holds since the algorithms immediately delete all further copies of an edge upon a deletion or recolouring. Indeed, suppose the graph $G_{v}$ is a graph with $|v|$ minimum that invalidates the upper-bound $\hat{\Delta}_{|v|}$. Then it must be because the colouring on $G_{|v|^{(|v|-1)}}$ is not a $t(|v|-1)$-splitter, or because $G_{v}$ at some point is not a subgraph of $G_{|v|^{(|v|-1)}}[v_{|v|}]$. The first case cannot happen, since the $t(|v|-1)$-splitter on $G_{|v|^{(|v|-1)}}$ always has all requirements satisfied by assumption. The second case cannot happen, since any time an edge coloured $v_{|v|}$ is deleted or recoloured in $G_{|v|^{(|v|-1)}}$, it is immediately removed from $G_{v}$. Hence, any time that $G_{v}$ is subjected to an insertion, it is a subgraph of $G_{|v|^{v}}[v_{|v|}]$ and thus satisfies the degree-bound.
Since we construct $\mathcal{G}_{i+1}$ by applying Theorem 3 to every graph in $\mathcal{G}_{i}$, and since Theorem 3 returns a valid $t(i)$-splitter, we thus also recover Item $4$. ∎

### v2

@@ 5 Maintaining a hierarchy of $t$-splitters
In this section, we will show how to build a hierarchy of dynamic $t$-splitters in order to split the degree all the way down to $O(\operatorname{\text{{\rm poly}}}(\varepsilon^{-1},\log n))$ while retaining an update time in $2^{\tilde{O}_{\log\varepsilon^{-1}}(\sqrt{\log n})}$. Thus we will show Theorem 1.4, which we restate below:
Theorem 5.1 (Identical to Theorem 1.4).
Let $G$ be a dynamic graph, and let $1>\varepsilon>0$ and an upper bound $\Delta_{\max}\geq\tfrac{10^{7}\log^{5}n}{\varepsilon^{3}}$ on the maximum degree of $G$ through the entire update sequence be given. Set $t_{1}=\lfloor 2^{10\sqrt{\log n}}\rfloor$ and $t_{2}=\lfloor\tfrac{\log n}{\varepsilon}\rfloor$. Then there exists a dynamic algorithm maintaining parameters $h,i,j$ with $h=i+j\leq 4\sqrt{\log n}$ and a set of graphs $\mathcal{G}_{h}$ satisfying that:
@@
  3. 3.
    $\hat{\Delta}_{i+j}\leq(1+\tfrac{\varepsilon}{16})\tfrac{\Delta_{\max}}{t_{1}^{i}t_{2}^{j}}\leq\operatorname{\text{{\rm poly}}}{\mathopen{}\left(\tfrac{\log n}{\varepsilon}\right)\mathclose{}}$
@@ 5.1 The hierarchy
We first describe the hierarchy in words. The hierarchy will consist of two types of $t$-splitters. The first type of $t$-splitter, we will use, has $t_{1}=\lfloor 2^{10\sqrt{\log n}}\rfloor$, and the second type has $t_{2}=\lfloor\tfrac{\log n}{\varepsilon}\rfloor$. Given any graph $G$, any $\Delta_{\max}$, and any $0<\varepsilon<1$, we will select some $h$. The hierarchy then consists of a sequence of sets of graphs $\{\mathcal{G}_{i}\}_{i=0}^{h}$ and a sequence of values $\{\hat{\Delta}_{i}\}_{i=0}^{h}$ such that for all $0\leq i\leq h$ and all $H\in\mathcal{G}_{i}$, we have that $\max_{v\in V(H)}d(v)\leq\hat{\Delta}_{i}$ i.e. the sequence of $\hat{\Delta}_{i}$.
@@
We note here that the algorithm from Theorem 1.3 is not history independent so ’running’ it on a static graph is ill-defined (one can think of the edges being processed in an arbitrary order). We will ignore this in the following informal description, and then provide a proper definition afterwards. We will initialise $\mathcal{G}_{0}=\{G\}$ and $\hat{\Delta}_{0}=\Delta_{\max}$, and proceed as follows: given $\mathcal{G}_{i}$ and $\hat{\Delta}_{i}$, we let $\mathcal{G}_{i+1}$ be the set of graphs obtained in the following manner. If $\Delta_{\max}\geq 2^{10i\sqrt{\log n}}$, we will run the algorithm from Theorem 1.3 with the parameter $\mu=\tfrac{\varepsilon^{\prime 2}}{\log n}$ and $t=t_{1}$ on all $H\in\mathcal{G}_{i}$. For each $H$, every colour-class $\kappa$ induces a subgraph $H_{\kappa}$ of $H$. For each choice of $H$ and $\kappa$, we include $H_{\kappa}$ in $\mathcal{G}_{i+1}$ after removing all degree 0 vertices. Finally, by Theorem 1.3, we can choose $\hat{\Delta}_{i+1}=(1+\mu)\tfrac{\hat{\Delta}_{i}}{t}$. Otherwise we set $t=t_{2}$ and repeat the above construction. We will pick $h$ such that the construction terminates with $\hat{\Delta}_{h}=O(\operatorname{\text{{\rm poly}}}(\varepsilon^{-1},\log n))$, but still large enough for any application of Theorem 1.3 to go through.
@@
The idea is that the edgesets of the graphs in $\mathcal{G}_{h}$ form a partition of the edges of $G$ such that each partition induces a low-degree subgraph of $G$. Since the $t$-splitters used are fully-dynamic, we can maintain the above hierarchy dynamically as well (see Algorithms 1 2 3). To do so, we simply build the $t$-splitters as specified, and whenever an edge is inserted, deleted or recoloured, we remove it from the graphs it previously belonged to (if any) and instead insert it into the graphs specified by its new colour (if any). The order of operations does not matter too much, as long as we guarantee that any $t$-splitter algorithm that is called is allowed to finish and that the $\hat{\Delta}_{i}$’s remain valid upper-bounds through-out the update sequence. In Algorithms 1 2 3 we will explicitly fix the order of operations.
@@
Next, we introduce some basic notation in order to make the above description formal. First of all, we define a $(G,\Delta_{\max},\varepsilon)$-valid-hierarchy. This is any hierarchy obtained by splitting via any $t$-splitter satisifying Invariant 4.2, and not necessarily only via the dynamic $t$-splitter from Theorem 1.3. This deals with the technicality that the data structure from Theorem 1.3 might not be history-independent, and so we cannot dynamically hope to maintain a very strictly defined hierarchy.
Definition 5.2.
@@
  • $\mathcal{G}_{i+1}:=\{\hat{H}:\exists H\in\mathcal{G}_{i},\kappa\in[t(i)]\text{ with }H[\kappa]=\hat{H}\text{ under }c_{H}\}$.
  • For all $i$: $\hat{\Delta}_{i+1}\leq(1+\tfrac{\varepsilon}{128\log n})\tfrac{\hat{\Delta}_{i}}{t(i)}$
@@
Remark 5.3.
We remark that the choice of the value $2^{(10+(i-1))\log n}$ is not extremely precise. One could have equally well chosen $2^{(9+(i-1))\log n}$ or $20\cdot{}2^{(10+(i-1))\log n}$. One then need only change the constant in Algorithms 1 3 2, in Definition 5.2, and in the subsequent proofs mutatis mutandi. In particular, one can always discretize the values and slightly over- or under-estimate all parameters like $\Delta_{\max},\varepsilon$ or $2^{(10+(i-1))\log n}$ to ensure that one does not need access to real RAM in order to do fast comparisons.
In particular, any hierarchy that uses a valid instantiation of Theorem 1.3 for every $t$-splitter yields a $(G,\Delta_{\max},\varepsilon)$-valid-hierarchy; regardless of the insertion order. We will also observe that for all $i$, the edge sets of the graphs in $\mathcal{G}_{i}$ form a valid partition of $E(G)$. We will let a vector $\chi(e)$ describe the hierarchy-profile of an edge:
Definition 5.4.
@@
For any $i\leq h$ and any vector $v\in\mathbb{Z}^{h}$, we let the $i^{\text{th}}$ initial segment of $v$ be the vector $v^{(i)}\in\mathbb{Z}^{i}$ whose first $i$ coordinates agree with the first $i$ coordinates of $v$. For any vector $v\in\mathbb{Z}^{i}$ and constant $k\in\mathbb{Z}$, we also let $v^{(i)},k]$ be the vector in $\mathbb{Z}^{i+1}$ with $i^{th}$ initial segment equal to $v$ and $v,k]_{i+1}=k$ We can index the graphs in the hierarchy by providing an index set $J$ containing every initial segment of the hierarchy-profile of every edge $e\in G$. Here, we let $G_{\emptyset}=G$, and we let $G_{j}$ for any $j\in J$ be the graph containing exactly the edges $e$ for which $j$ is an initial segment of $\chi(e)$. We let $|j|:=dim(j)$ denote the dimension of the vector $j$. Also $|j|$ is the level of $G_{j}$.
@@ 5.2 An algorithm for maintaining a hierarchy
The algorithm for dynamically maintaining the hierarchy is quite straight-forward. First of all, one initializes values $t(i)$ and $\hat{\Delta}_{i}$ so that they conform to the constraints of Definition 5.2 (see Algorithm 1). This only has to be done once.
The hierarchy is then maintained by repeatedly applying Theorem 1.3 to further split graphs into ones with lower degree. First, one performs the update on level 0, which then propagates the update along with any extra recolourings to the next level of the hierarchy. One has to be a bit careful with determining the order of deletions and insertions, in order to ensure that the $\hat{\Delta}_{i}$’s remain valid upper-bounds through-out the update sequence. The algorithm will delete in a bottom-up fashion, beginning at the bottom of the hierarchy, and insert in a top-down fashion. That is every time an edge is deleted or recoloured, the algorithm first deletes every affected instance of the edge, beginning at the bottom. When an edge is inserted (or receives a new colour), the edge is inserted into the top levels of the structure first, in order to ensure that it does not enter a lower level structure before it is guaranteed that the maximum degree is preserved. After processing an insertion or deletion at a level some edges at the level are (potentially) recoloured. In order to process the recoloured edges, we ensure to first perform all deletions, and subsequently the insertions. Thus, we can ensure that the degree bounds hold at all times before an insertion in to a graph is processed. This ensures that $\hat{\Delta}_{i}$’s remain valid upper-bounds after the updates. Finally, we note that the hierarchy consists of many graphs, which can either be constructed in a pre-processing phase or simply allocated online by only inserting vertices into a graph once the vertex actually has degree at least one.
Next, we give pseudo-code for the above dynamic algorithm. If an edge $e$ has just been recoloured, we let $c(e)_{\text{old}}$ respectively $\chi(xy)_{\text{old}}$ denote the colours of $e$ just before the recolouring. The hierarchy is maintained via the following set of recursive procedures. The first procedure shows how to initialize the parameters needed for the following procedures. We note already now that to actually perform an insertion or a deletion, one must call the respective recursive procedure with the correct parameter setting (see Remark 5.5):
@@
$\mu=\lfloor\tfrac{\varepsilon}{128\log n}\rfloor$
@@
   $\hat{\Delta}_{i_{1}+i_{2}}=(1+\mu)^{i_{1}+i_{2}}\tfrac{\Delta_{\max}}{t_{1}^{i_{1}}t_{2}^{i_{2}}}$
@@
   $\hat{\Delta}_{i_{1}+i_{2}}=(1+\mu)^{i_{1}+i_{2}}\tfrac{\Delta_{\max}}{t_{1}^{i_{1}}t_{2}^{i_{2}}}$
  end if
end while
@@
[…]
insert(uv) into $G_{j}$ using Theorem 1.3 with $(\Delta_{\max},t,\varepsilon^{\prime})=(\hat{\Delta}_{|j|},t(|j|),\mu)$
@@
end for
[…]
Algorithm 2 Insert-Hierarchy(uv,j)(uv,j)
delete(uv) from $G_{j}$ using Theorem 1.3 with $(\Delta_{\max},t,\varepsilon^{\prime})=(\hat{\Delta}_{|j|},t(|j|),\mu)$
if $|\zeta|\geq|j|+1$ then
[…]
for all recoloured edges $xy\neq uv$ in $G_{j}$ do
[…]
end for
[…]
Remark 5.5.
We remark that to actually perform an insertion, one calls $\texttt{Insert-Hierarchy}(uv,\emptyset)$, and to actually perform a deletion, one calls $\texttt{Delete-Hierarchy}(uv,\emptyset,\chi(uv))$.
5.3 Correctness
In this section, we show that the algorithms from above actually maintain a $(G,\Delta_{\max},\varepsilon)$-valid-hierarchy with parameters that give the guarantees stated in Theorem 1.4. We first show that the algorithm maintains a $(G,\Delta_{\max},\varepsilon)$-valid-hierarchy:
Lemma 5.6.
Algorithms 1 2 3 implement a $(G,\Delta_{\max},\varepsilon)$-valid-hierarchy.
Proof 13.
The proof is a straight-forward verification that the hierarchy returned satisfies the items in Definition 5.2. To that end, we note that $\texttt{Initialize-Parameters}(\Delta_{\max},\varepsilon)$ initialises in accordance with item 1. This part is never changed, and so through-out all updates Item 1 holds.
Furthermore, $\texttt{Initialize-Parameters}(\Delta_{\max},\varepsilon)$ ensures that the sequence $\{t(i)\}_{i=0}^{h-1}$ is in accordance with the second item of Definition 5.2.
The hierarchy is also in accordance with the third, fifth, and sixth items. Indeed, if at all times the conditions of Theorem 1.3 is met, then it follows by induction and the theorem itself that the computed colourings are $t(i)$-splitting for all $i$, and furthermore satisfy the fifth and sixth items.
To check that this is the case, we check that the upper bounds on the degrees $\hat{\Delta}_{i}$ remain valid at all times during an update. This holds since the algorithms immediately delete all further copies of an edge upon a deletion or recolouring. Indeed, suppose the graph $G_{v}$ is a graph with $|v|$ minimum that invalidates the upper-bound $\hat{\Delta}_{|v|}$. Then it must be because the colouring on $G_{v^{(|v|-1)}}$ is not a t(|v|1)t(|v|-1)-splitter, or because $G_{v}$ at some point is not a subgraph of $G_{v^{(|v|-1)}}[v_{|v|}]$. The first case cannot happen, since the t(|v|1)t(|v|-1)-splitter on $G_{v^{(|v|-1)}}$ always has all requirements satisfied by assumption. The second case cannot happen, since any time an edge coloured $v_{|v|}$ is deleted or recoloured in $G_{v^{(|v|-1)}}$, it is immediately removed from $G_{v}$. Hence, any time that $G_{v}$ is subjected to an insertion, it is a subgraph of $G_{v^{(|v|-1)}}[v_{|v|}]$ and thus satisfies the degree-bound.
Since we construct $\mathcal{G}_{i+1}$ by applying Theorem 1.3 to every graph in $\mathcal{G}_{i}$, and since Theorem 1.3 returns a valid $t(i)$-splitter, we thus also recover Item 44.